# Analytical study of a finite-range impurity in a one-dimensional Bose gas

### v1

@@ IV Universality in the Strongly Coupled Regime
$x_{r}=\frac{\coth^{2}(z_{r})-1}{2z_{r}^{2}}\left(\frac{\tilde{w}}{2}\right)^{3}.$ (12)
Once $x_{r}$ is determined, the energy of the repulsive polaron can be approximated by evaluating the integral (6), assuming $\phi_{0}(x)=\tanh((|x|+x_{r})/(\sqrt{2}\xi))$ holds everywhere:
@@
We derive an identity, valid for all $V_{0}$, specifically useful in determining the phase function $\phi_{1}(x)$ in the region $|x|\leq w/2$. One can obtain the expression $\partial_{xx}\phi_{0}/\phi_{0}=\phi_{0}^{2}-1+U_{IB}(x)$ from (4). Substituting this into (5) and integrating by parts over the domain $x\in[0,\infty)$ yields:
@@
To compute the term $\partial_{x}\phi_{1}({w}/2)$, we note that $\phi_{0}(|x|<w/2)\simeq\sqrt{2W_{r}\epsilon_{r}}\cosh(\sqrt{W_{r}}x)+\mathcal{O}(\epsilon_{r}^{3/2})$. Therefore, the solution of Eq. (5) is
$\phi_{1}\left(|x|<w/2\right)\simeq-\frac{\sinh(z_{r})}{\sqrt{2}}\sinh(\sqrt{W_{r}}x)+\mathcal{O}(\epsilon_{r}^{3/2}).$ (16)
@@
For the attractive impurity in the strong coupling limit, we have $\tilde{V_{0}}\simeq\left(\frac{z_{a}}{\tilde{w}/2}\right)^{2}$, where $z_{a}=-\cot z_{\alpha}$ (see Eq. (23)). The solution (Eq. (8)) approaches the form
$\phi_{0}(x)\simeq\sqrt{W_{a}}\left[1-\frac{\epsilon_{a}}{4}(\cosh(\sqrt{\frac{W_{a}}{2}}x)-1)\right]$ (18)
@@
$x_{a}=\frac{\tilde{w}}{2}\left(\frac{\sqrt{2}}{z_{a}}\frac{1}{1+\frac{\epsilon_{a}}{4}\left(1-\cosh(\sqrt{2}z_{a})\right)}-1\right).$ (20)
@@ V Conclusion
We derived an analytical solution to the ground-state wave function of the GP equation for a one-dimensional Bose gas with a finite-ranged impurity and calculated perturbative corrections arising from the impurity’s slow motion. This approach resolves the divergence of polaron properties in the strongly coupled regime observed with contact potentials. Moreover, we found that in this regime the polaron properties admit a universal description determined by the ratio of the interaction range to the coherence length. The developed formalism can also be applied to study polarons in the long-range regime, $\xi\gg w$, and to obtain excited-state solutions of the GPE, enabling investigation of the connection between polaron branches and GPE solutions in one dimension.
@@ Appendix A Scattering Length of the Finite-Well Impurity
Here, we briefly summarize the bound and scattering states of the square well potential. We restrict our analysis to the even-parity solution, since the ground-state wavefunction exhibits even parity.
The relation between the scattering length $a_{IB}$ and the parameters of the potential depth is:
[…]
for the attractive and repulsive finite-well potential [36, 37], respectively. Here $z=\sqrt{\tilde{V}_{0}}\frac{w}{2\tilde{\xi}}$ and $\tilde{V}_{0}=V_{0}/(gn_{0})$. In contrast to the three-dimensional case, where the coupling strength $g_{IB}$ is proportional to the scattering length $a_{IB}$, the relationship in one dimension is both sign-reversed and inverted, given by $g_{IB}=-\frac{\hbar^{2}}{m_{r}a_{IB}}$. We illustrate the mapping between the square-well parameters, and the scattering length $a_{IB}$ and the coupling strength $g_{IB}$ in Fig. 5.
@@
An infinitesimally shallow attractive square well always supports an even-parity bound state. Additional odd- and even-parity states appear as $z$ crosses $(n-\tfrac{1}{2})\pi$ and $n\pi$, respectively, with $n$ a positive integer. The emergence of a new even-parity state coincides with a scattering-length divergence from -$\infty$ to $\infty$. A repulsive square well supports no bound states and thus no such transition. Notably, the strongly coupled limit ($g_{IB}\to\pm\infty$) does not coincide with the unitarity points where new bound states form.
@@
Figure 5: The mapping between the parameters $z=\sqrt{V_{0}}\frac{w}{2}$ of the square-well potential and the corresponding zero-energy scattering length $a_{IB}$ (blue lines), as well as the coupling constant $g_{IB}=-\frac{\hbar^{2}}{m_{r}a_{IB}}$ (red lines), is shown for (a) Attractive, and (b) Repulsive impurity.
@@
We model the repulsive impurity using a repulsive square-well potential ($V_{0}>0$) as a simplified representation of the impurity–bath interaction, interpreting the repulsive polaron as the ground-state solution under this potential. While this approach does not capture the metastable nature of repulsive polarons - typically arising from attractive short-range interactions- it remains valuable for gaining physical insights into their behavior and exploring how their properties depend on the parameters of a finite-range potential. Analogous use of repulsive potentials to approximate short-range interactions between bath particles has yielded physically consistent results [38].
Appendix B Derivation of the GP Equation
The Hamiltonian of an impurity-bath system with the chemical potential $\mu$ is represented in the second-quantized formalism as:
$+$ $\hat{U}_{IB}(x-\hat{x}_{I})-\mu\bigg{]}\hat{\Psi}(x),$
[…]
$\hat{\mathcal{K}}_{LLP}=\frac{(p_{\mathrm{tot}}-\hat{p}_{B})^{2}}{2m_{I}}+\int dx\hat{\Psi}^{\dagger}(x)\bigg{[}-\frac{\hbar^{2}\hat{\partial}_{x}^{2}}{2m_{r}}+\frac{g}{2}\hat{\Psi}^{\dagger}(x)\hat{\Psi}(x)+\hat{U}_{IB}(x)-\mu\bigg{]}\hat{\Psi}(x),$ (26)
where $m_{r}$ is the reduced mass. Note that the Hamiltonian (26) is independent of the impurity’s position $x_{I}$ and explicitly incorporates the conserved total momentum $p_{\mathrm{tot}}\mathbf{e_{x}}$ of the impurity–bath system. We proceed by applying the mean-field (MF) approximation to the Hamiltonian in (26), replacing the operators by their expectation values, $\hat{\Psi}(x)\simeq\langle\hat{\Psi}(x)\rangle=\Psi(x)$, while neglecting any fluctuations. We apply the variational analysis $\frac{\delta\mathcal{K}_{LLP}}{\delta\Psi^{*}}=0$ to derive the time-independent GPE governing the Bose field with non-zero total system momentum $p_{\mathrm{tot}}$:
$\mu\Psi(x)$ $=$ $-\frac{\hbar^{2}\partial_{x}^{2}\Psi(x)}{2m_{r}}+U_{IB}(x)\Psi(x)+g|\Psi(x)|^{2}\Psi(x)$
where the term in square brackets in the second line is the impurity momentum $p_{I}=p_{\mathrm{tot}}-p_{B}$ at the MF level, with $\hat{p}_{B}=-i\hbar\int dx\hat{\Psi}^{\dagger}(x)\partial_{x}\hat{\Psi}(x)$ as the total phonon momentum. We set the chemical potential $\mu=gn_{0}$, where $n_{0}$ is the uniform bath density far from the impurity.
Appendix C Properties of Jacobi elliptic functions
@@
Appendix D The Equations of Boundary Conditions

### v2

@@ IV Polaron Properties in the Strongly Coupled Regime
$\frac{x_{r}}{\tilde{\xi}}=\frac{\coth^{2}(z_{r})-1}{2z_{r}^{2}}\left(\frac{\tilde{w}}{2}\right)^{3}.$ (17)
Once $x_{r}$ is determined, the energy of the repulsive polaron can be approximated by evaluating the integral (11), assuming $)\phi_{0}(x)=\tanh((|x|+x_{r})/(\sqrt{2}\tilde{\xi}$)) holds everywhere:
@@
We derive an identity, valid for all $V_{0}$, specifically useful in determining the phase function $\phi_{1}(x)$ in the region $|x|\leq w/2$. One can obtain the expression $\partial_{xx}\phi_{0}/\phi_{0}=\phi_{0}^{2}-1+\tilde{U}_{\mathrm{IB}}(x)$ from (9). Substituting this into (10) and integrating by parts over the domain $x\in[0,\infty)$ yields:
@@
To compute the term $\partial_{x}\phi_{1}({w}/2)$, we note that $\phi_{0}(|x|<w/2)\simeq\sqrt{2W_{r}\epsilon_{r}}\cosh(\sqrt{W_{r}}x/\tilde{\xi})+\mathcal{O}(\epsilon_{r}^{3/2})$. Therefore, the solution of Eq. (10) is
$\phi_{1}\left(|x|<w/2\right)\simeq-\frac{\sinh(z_{r})}{\sqrt{2}}\sinh(\sqrt{W_{r}}x/\tilde{\xi})+\mathcal{O}(\epsilon_{r}^{3/2}).$ (21)
@@
For the attractive impurity in the strong coupling limit, we have $\tilde{V_{0}}\simeq\left(\frac{z_{a}}{\tilde{w}/2}\right)^{2}$, where $z_{a}=-\cot z_{a}\simeq 2.798$ (see Eq. (6)). The solution (Eq. (13)) approaches the form
$\phi_{0}(x)\simeq\sqrt{W_{a}}\left[1-\frac{\epsilon_{a}}{4}(\cosh(\sqrt{\frac{W_{a}}{2}}\frac{x}{\tilde{\xi}})-1)\right]$ (23)
@@
$\frac{x_{a}}{\xi}=\frac{\tilde{w}}{2}\left(\frac{\sqrt{2}}{z_{a}}\frac{1}{1+\frac{\epsilon_{a}}{4}\left(1-\cosh(\sqrt{2}z_{a})\right)}-1\right).$ (25)
@@ V Conclusion
We derived an analytical solution to the ground-state wave function of the GP equation for a one-dimensional Bose gas with a finite-ranged impurity and calculated perturbative corrections arising from the impurity’s slow motion. This approach resolves the divergence of polaron properties in the strongly coupled regime observed with contact potentials. Moreover, we found that in this regime the polaron properties admit a description determined by the ratio of the interaction range to the coherence length. The developed formalism can also be applied to study polarons in the long-range regime, $\xi\gg w$, and to obtain excited-state solutions of the GP equation, enabling investigation of the connection between polaron branches and GP equation solutions in one dimension.
@@ Appendix A Scattering Length of the Finite-Well Impurity
Here, we briefly summarize the bound and scattering states of the square well potential. We restrict our analysis to the even-parity solution, since the ground-state wavefunction exhibits even parity. We illustrate the mapping between the square-well parameters, and the scattering length $a_{\mathrm{IB}}$, Eq. (6) and Eq. (7), and the coupling strength $g_{\mathrm{IB}}$ in Fig. 5.
@@
An infinitesimally shallow attractive square well always supports an even-parity bound state. Additional odd- and even-parity states appear as $z$ crosses $(n-\tfrac{1}{2})\pi$ and $n\pi$, respectively, with $n$ a positive integer. The emergence of a new even-parity state coincides with a scattering-length divergence from -$\infty$ to $\infty$. A repulsive square well supports no bound states and thus no such transition. Notably, the strongly coupled limit ($g_{\mathrm{IB}}\to\pm\infty$) does not coincide with the unitarity points where new bound states form.
@@
Figure 5: The mapping between the parameters $z=\sqrt{\frac{|V_{0}|2m_{r}}{\hbar^{2}}}\frac{w}{2}$ of the square-well potential and the corresponding zero-energy scattering length $a_{\mathrm{IB}}$ (blue lines), as well as the coupling constant $g_{\mathrm{IB}}=-\frac{\hbar^{2}}{m_{r}a_{\mathrm{IB}}}$ (red lines), is shown for (a) attractive, and (b) repulsive impurity. The gas parameter $\gamma=1/(2n_{0}^{2}\xi^{2})=0.438$, and the impurity-bath mass ratio $m_{I}/m=0.47$ are used to obtain the coupling strength $\eta$.
@@
We model the repulsive impurity using a repulsive square-well potential ($V_{0}>0$) as a simplified representation of the impurity–bath interaction, interpreting the repulsive polaron as the ground-state solution under this potential. While this approach does not capture the metastable nature of repulsive polarons - typically decay into the ground state arising from attractive short-range interactions- it remains valuable for gaining physical insights into their behavior and exploring how their properties depend on the parameters of a finite-range potential. Analogous use of repulsive potentials to approximate short-range interactions between bath particles has yielded physically consistent results [53].
Appendix B Properties of Jacobi elliptic functions
@@
Appendix C The Equations of Boundary Conditions